Let x be the solution of the equation cosec(x)=3 range:pi/2
Answer by stanbon(75887)

Let x be the solution of the equation cosec(x)=3
range:pi/2 If csc(x) = 3, sin(x) = 1/3 in QII
-------------
x = sin^-1(1/3) in QII = pi-0.34 = 2.80 radians
-------
Then x=pi- arctan (y) what is y?
2.80 = pi - tan^-1(y)
tan^-1(y) = pi - 2.80
tan^-1(y) = 0.34
y = 0.3277
================
Cheers,
Stan H.